# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v1

@@ 6.3 Evaluation Metrics for Agentic AI Systems
Agentic AI systems demand a comprehensive evaluation across multiple dimensions beyond traditional accuracy. We outline five key categories of metrics: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct aspect of an Agentic AI’s performance and its real-world implications.. Below we discuss these metrics and summarize in Table 7.
Trustworthiness: This category evaluates the reliability, safety, and ethical alignment of the AI agent. A trustworthy agent consistently produces correct and unbiased results, adheres to constraints, and avoids harmful or unpredictable behavior. Metrics for trustworthiness include success rates on tasks under varying conditions (measuring robustness), violation rates of safety or ethical guidelines (which should be minimal), and calibration of the agent’s confidence (how well the agent’s self-reported confidence aligns with actual accuracy) liu2023gevalnlgevaluationusing . Some approaches combine such factors into an overall trust score. For example, one model defines a trustworthiness score $T$ as
@@
Explainability: Explainability metrics assess how well the agent’s decisions can be understood and traced by humans munoz2024explainability . These metrics focus on the clarity and completeness of the rationale the agent provides for its actions. For instance, one can measure the explanation coverage (the percentage of decisions or outputs that come with an adequate explanation) and the fidelity of explanations (how accurately the explanation reflects the true reasoning or model logic).
Consistency of explanations for similar scenarios is another important metric: the agent should explain comparable decisions in a similar way, indicating a stable reasoning process. Quantitatively, methods like OpenXAI provide a suite of metrics to evaluate explanation quality across dimensions such as faithfulness, stability, and fairness of explanations agarwal2024openxaitransparentevaluationmodel . High explainability builds user trust, as users can follow why the agent made a decision, and it aids debugging by revealing the agent’s internal decision process. In regulated domains (e.g., healthcare or finance), explainability is often essential for compliance and user acceptance.
User-Centered Performance: User-centered metrics capture how effectively the AI agent interacts with and satisfies the end-user’s needs. These criteria emphasize the user’s experience and outcomes rodden2010measuring . Key metrics include user satisfaction ratings, typically collected via surveys or feedback after interaction, which reflect whether the user’s goals were met and how comfortable they felt with the agent’s behavior. Task success from the user’s perspective (did the agent fulfill the user’s request or solve the user’s problem?) is a fundamental measure.
Additionally, interaction metrics like the number of back-and-forth clarification queries needed (fewer indicates the agent understood the user well) and the coherence or naturalness of the conversational flow contribute to user-centered evaluation. Human-in-the-loop evaluations are often employed here: for example, user studies might rate the agent on criteria such as helpfulness, clarity and naturalness of language, and adherence to user instructions. Ultimately, a user-centered agentic system should align its actions with user intent and preferences.
Benchmarks like ChatDev qian2023chatdev , which simulates a multi-agent software development team interacting via natural language, implicitly evaluate how well agents fulfill user-defined roles and requirements in a collaborative project. This highlights the importance of user-oriented success in complex, realistic tasks.
@@
Coordination (Multi-Agent or Modular): In scenarios where an Agentic AI system consists of multiple cooperating agents or modular components, coordination metrics gauge how effectively these parts work together. Good coordination means agents share information, divide labor without conflict or redundancy, and converge on solutions efficiently. Quantitative measures include the team success rate on collaborative tasks (whether the group of agents achieves the overall goal) and communication efficiency metrics (e.g., the number of messages or iterations required among agents to reach a decision, with fewer often indicating more efficient interaction).
One specific example is the Component Synergy Score (CSS), which counts or weights effective interactions between agents, reflecting how well each agent’s actions complement the others (a higher CSS means agents are synergistic rather than working at cross-purposes). Multi-agent frameworks such as ChatDev and MetaGPT provide practical testbeds for these metrics: they orchestrate specialized agents (e.g., different roles in a software engineering pipeline) that must cooperate to complete complex projects.
@@
Composite Metrics: Composite metrics aggregate multiple evaluation aspects into a single overall score. These are useful for summarizing an agent’s performance holistically, especially when comparing different systems. A composite metric is often a weighted combination of the above categories, for example:
@@
where $M_{T},M_{E},M_{U},M_{C}$ are the normalized scores (on a common scale) for trustworthiness, explainability, user-centered performance, and coordination respectively, and $w_{T},w_{E},w_{U},w_{C}$ are weights reflecting the relative importance of each aspect for a given application. The choice of weights $w_{i}$ can be domain-specific (for instance, in healthcare applications, trustworthiness and explainability might be weighted more heavily than raw efficiency). An example of a specialized composite metric in agentic contexts is the Tool Utilization Efficacy (TUE) score, which combines sub-metrics evaluating how correctly and efficiently an agent uses external tools (including proper tool selection and correct parameter usage in tool calls) into one measure. By condensing multiple criteria, composite metrics enable high-level comparison and benchmarking of agentic systems. For instance, AgentBench liu2023agentbench is a comprehensive benchmark that evaluates agents across a diverse range of tasks and environments (from operating system manipulation to web shopping), effectively providing a composite performance profile of an agent. Such aggregated scores highlight if an agent performs strongly across the board or if it excels in some dimensions while underperforming in others. It is important to interpret composite scores in light of their components: a single number can mask specific weaknesses (e.g., an agent might achieve a high overall score by doing well in task completion and coordination, yet still have poor explainability). Therefore, composite metrics are most informative when accompanied by a breakdown of the agent’s performance per category.
@@ 7.1 Security Mechanisms
Agentic AI systems, composed of loosely coupled yet collaboratively functioning LLM-based agents, introduce an expanded attack surface relative to conventional AI agents khan2024security . Ensuring the security of such systems necessitates a multi-layered defense architecture that addresses data protection, execution integrity, inter-agent communication, and model robustness wang2024large . Among the foundational techniques employed are encryption he2024security ; motwani2024secret , access control huang2025agent ; south2025authenticated , adversarial defense shayea2025strategies ; malik2024systematic , and runtime monitoringnarajala2025securing ; chan2024agentmonitor each adapted to the unique demands of decentralized multi-agent environments.
@@
Encryption plays an important role in safeguarding data exchanged between agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved motwani2024secret ; feretzakis2024privacy ; bommi2025enhancing . Agentic workflows often include inter-agent handoff of partially processed results, models, or prompts. Implementations such as SSL/TLS, homomorphic encryption jency2025finquaxbot ; bommi2025enhancing , and secure enclaves (e.g., Intel SGX  [^1] https://www.intel.com/content/www/us/en/products/docs/accelerator-engines/software-guard-extensions.html) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
@@
Access control becomes crucial when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities. For instance, in systems like AutoGenwu2023autogen and CrewAI crewAI where agents take on specialized roles (e.g., summarizer, planner, coder) epperson2025interactive , enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation wu2023autogen ; narajala2025securing ; ismail2025toward . Agent-based access control policies often aligned with Role-Based Access Control (RBAC) khan2024role ; huang2025novel and Attribute-Based Access Control (ABAC) neupane2025towards ; huang2025novel paradigms can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels sapkota2025ai .
@@
Adversarial robustness is a growing concern as LLM-based agents are susceptible to prompt injection, manipulation through poisoned tool outputs, or coordination disruption via malformed intermediate results liu2024exploring ; zhang2024breaking . Recent studies have shown that multi-agent LLM frameworks can be destabilized by adversarially crafted outputs from one compromised agent propagating misleading information to others zhang2024breaking ; wang2024large ; deng2025ai ; wang2025comprehensive ; patlan2025real ; patlan2025ai . Adversarial training methods, such as input perturbation, reward shaping, and contrastive learning, can partially mitigate these vulnerabilities standen2025adversarial ; deng2025ai . Integrating safety constraints and verifying tool responses before execution are also effective mitigation strategies.
@@
Runtime monitoring systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity chan2024visibility . Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers zhen2025novel ; sapkota2025vibe ; xu2025forewarned . For example, Microsoft’s Copilot governance layers monitor anomalous agent behavior across sessions to ensure compliant execution and flag potentially harmful interactions.
@@
As Agentic AI continues to expand into mission-critical domains, developing standardized, scalable security mechanisms will be paramount. Future approaches must include zero-trust frameworks, secure multi-party computation, and formal verification of inter-agent protocols to ensure safe and resilient operation across decentralized autonomous agent collectives.
@@ 7.2 Privacy-Preserving Techniques
Differential privacy (DP) offers mathematically grounded guarantees by injecting statistical noise into outputs, ensuring that individual user contributions cannot be re-identified. In multi-agent LLM systems, DP can be applied during training or at inference-time when agents exchange information. For example, Google’s implementation of DP in federated learning frameworks can be extended to distributed agentic systems, where agents collaboratively train or fine-tune local models without exposing raw data lazaros2024federated ; luzon2024tutorial . DP-SGD and privacy budgets ($\epsilon$-differentials) can regulate information exposure during policy updates or collaborative planning in real-time decision-making agents.
Data minimization is another cornerstone of privacy preservation. Agentic AI systems can mitigate exposure risks by limiting the scope, granularity, and duration of data collected or retained during task execution. For instance, temporary memory buffers used in systems like ChatDev or ReAct-based pipelines are cleared once subgoals are completed, preventing persistent storage of unnecessary user data. Furthermore, anonymization and pseudony luzon2024tutorial , techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
Secure computation techniques including secure multi-party computation (SMPC), homomorphic encryption, and trusted execution environments (TEEs) enable agents to perform computations over encrypted or obfuscated data without compromising privacy. In scenarios where agents collaborate across different organizational boundaries (e.g., federated medical agents or cross-silo industrial agents), SMPC allows joint computations such as diagnostics or anomaly detection without data leakage. Homomorphic encryption, while computationally expensive, is increasingly being explored to allow arithmetic operations on encrypted vectors used in RAG workflows.
@@
Table 8: Governance Dimensions for Agentic AI Systems
Governance Component Description Example Tools or Methods
Regulatory Compliance Aligning with legal and industry standards for responsible AI raza2025responsible NIST AI RMF dotan2024evolving , EU AI Act euregulationai2024 , GDPR gdpr25 , HIPAA hipaa164
Auditability and Logging Capturing traceable records of agent actions and decisions south2025authenticated Decision provenance, immutable logs, blockchain-based traceability
Policy Enforcement Enforcing system-level operational, ethical, and security constraints de2025open ; chen2024trustworthy RBAC/ABAC policies aba2023ai , formal logic rules, dynamic sandboxing zhong2023towards ; ali2018agent ; golechha2025among
Human Oversight Human-in-the-loop chen2025human decision checkpoints and override capabilities Human review agents, interactive dashboards, compliance checkpoints
Risk Monitoring Continuous identification and mitigation of systemic or emergent risks slattery2025airiskrepositorycomprehensive Model risk scanners, out-of-distribution detectors, reinforcement learning audit modules
Explainability Governance Ensuring interpretability standards are met in decision pipelines arrieta2020explainable LIME/SHAP salih2025perspective ; donmez2025explainable ; knab2025lime , decision trees yamada2025ai ; walia2025computational , interpretable surrogate models fonseca2025optimal ; mahmud2025optimizing ; xi2025model
Adaptive Governance Updating governance mechanisms based on model evolution and deployment context kolt2025governing ; hughes2025ai Governance-as-code Responsi73:online ; kose2025modern , auto-updated rule engines goh2024campus ; yu2024shadow ; yun2025generative , learning-based policy adaptation feng2025multi ; li2025investigation
@@ 7.3 Compliance and Governance
As Agentic AI systems grow in capability and autonomy, ensuring regulatory compliance and instituting robust governance mechanisms become imperative. Unlike traditional AI agents, agentic systems operate with greater autonomy, persistent memory, and complex decision flows necessitating layered oversight to manage legal, ethical, and societal implications. Effective governance in this context spans three critical dimensions: adherence to regulatory standards, system-level auditability, and enforceable policy frameworks.
@@
Regulatory standards provide the baseline requirements that all AI systems, including agentic architectures, must meet. Frameworks such as the NIST AI Risk Management Framework (AI RMF) and the EU AI Act define principles for trustworthy AI including transparency, accountability, and fairness. These standards are especially relevant for LLM-based AMAS that interact in high-stakes domains like healthcare, finance, defense, or transportation. For example, the EU AI Act classifies certain autonomous systems as “high-risk,” requiring continuous risk monitoring, documentation of decision logic, and human oversight mechanisms attributes directly relevant to Agentic AI.
@@
Auditability is critical in ensuring transparency and facilitating post-hoc accountability. Each decision, plan, or interaction within an agentic system should be logged with timestamps, context, agent role, and justification. Techniques such as decision provenance and action traceability enable this, allowing regulators or internal auditors to reconstruct how decisions were reached. For instance, in systems like AutoGen or MetaGPT, where agents assume specialized roles (e.g., researcher, coder, reviewer), audit trails can capture role-specific actions and flag inconsistencies, bias amplification, or security violations. Blockchain-based audit logs are also being explored to ensure immutability and verifiability of multi-agent interactions.
@@
Policy enforcement governs what agentic systems can and cannot do, and under which conditions. These policies must be codified into the orchestration layers or meta-agent governance modules that manage agent interactions. Examples include enforcing memory expiration policies to avoid data retention violations or restricting access to external tools based on role and authentication level. Role-based access control (RBAC) and attribute-based access control (ABAC) are essential for enforcing differentiated privileges across agent subcomponents. Furthermore, real-time monitoring systems can halt or flag agent activity that deviates from pre-specified ethical constraints or operational bounds, using formal verification tools like TLA+ or symbolic execution engines.
@@
Emerging best practices also include the creation of AI governance boards, the adoption of governance-as-code platforms, and the integration of adaptive governance layers that evolve as the agentic system scales or changes context. These practices are designed to meet not only current standards like ISO/IEC 42001 for AI management systems, but also prepare for future regulatory evolutions. Below, we summarize the core governance components required for secure and compliant Agentic AI deployment.
Thus, governance in Agentic AI systems is not a static compliance checkbox but a dynamic, adaptive layer embedded into the orchestration and operational pipelines. As these systems become more autonomous and integrated across sectors, aligning governance with system behavior, human values, and evolving legal standards is imperative for safe and trustworthy deployment, as summarized in Table 8.

### v2

@@ 7.3 Evaluation Metrics for Agentic AI Systems
Agentic AI systems demand a comprehensive evaluation across multiple dimensions beyond traditional accuracy. We outline five key categories of metrics: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct aspect of an Agentic AI system performance and its real-world implications. Below, we discuss these metrics and summarize in Table 7.
Trustworthiness. This criteria evaluates the reliability, safety, and ethical alignment of the AI agent. A trustworthy agent consistently produces correct and unbiased results, adheres to constraints, and avoids harmful or unpredictable behavior. Metrics for trustworthiness include success rates on tasks under varying conditions (measuring robustness), violation rates of safety or ethical guidelines (which should be minimal), and calibration of the agent’s confidence (how well the agent’s self-reported confidence aligns with actual accuracy) [164]. Some approaches combine such factors into an overall trust score. For example, one model defines a trustworthiness score $T$ as
@@
Explainability. Explainability metrics assess how well the agent’s decisions can be understood and traced by humans [165]. These metrics focus on the clarity and completeness of the rationale the agent provides for its actions. For instance, one can measure the explanation coverage (the percentage of decisions or outputs that come with an adequate explanation) and the fidelity of explanations (how accurately the explanation reflects the true reasoning or model logic).
Consistency of explanations for similar scenarios is another important metric: the agent should explain comparable decisions in a similar way, indicating a stable reasoning process. Quantitatively, methods like OpenXAI [166] provide a suite of metrics to evaluate explanation quality across dimensions such as faithfulness, stability, and fairness of explanations. High explainability builds user trust, as users can follow why the agent made a decision, and it aids debugging by revealing the agent’s internal decision process. In regulated domains (e.g., healthcare or finance), explainability is often essential for compliance and user acceptance.
User-Centered Performance. User-centered metrics capture how effectively the AI agent interacts with and satisfies the end-user’s needs. These criteria emphasize the user’s experience and outcomes [167]. Key metrics include user satisfaction ratings, typically collected via surveys or feedback after interaction, which reflect whether the user’s goals were met and how comfortable they felt with the Agentic AI system behavior.
Additionally, interaction metrics like the number of back-and-forth clarification queries needed (fewer indicates the agent understood the user well) and the coherence or naturalness of the conversational flow contribute to user-centered evaluation. Human-in-the-loop evaluations are often employed here: for example, user studies might rate the agent on criteria such as helpfulness, clarity and naturalness of language, and adherence to user instructions. Ultimately, a user-centered Agentic AI system should align its actions with user intent and preferences.
@@
Coordination. In scenarios where an Agentic AI system consists of multiple cooperating agents or modular components, coordination metrics gauge how effectively these parts work together. Good coordination means agents share information, divide labor without conflict or redundancy, and converge on solutions efficiently. Quantitative measures include the team success rate on collaborative tasks (whether the group of agents achieves the overall goal) and communication efficiency metrics (e.g., the number of messages or iterations required among agents to reach a decision, with fewer often indicating more efficient interaction).
One specific example is the Component Synergy Score (CSS), which counts or weights effective interactions between agents, reflecting how well each agent’s actions complement the others (a higher CSS means agents are synergistic rather than working on individual basis). MAS such as ChatDev and MetaGPT provide practical testbeds for these metrics: they orchestrate specialized agents (e.g., different roles in a software engineering pipeline) that must cooperate to complete complex projects.
@@
Composite Metrics. Composite metrics aggregate multiple evaluation aspects into a single overall score. These are useful for summarizing agents’ performance holistically, especially when comparing different systems. A composite metric is often a weighted combination of the above categories, for example:
@@
where $M_{T},M_{E},M_{U},M_{C}$ are the normalized scores (on a common scale) for trustworthiness, explainability, user-centered performance, and coordination, respectively, and $w_{T},w_{E},w_{U},w_{C}$ are weights reflecting the relative importance of each aspect for a given application. The choice of weights $w_{i}$ can be domain-specific (for instance, in healthcare applications, trustworthiness and explainability might be weighted more heavily than raw efficiency).
An example of a specialized composite metric in agentic contexts is the Tool Utilization Efficacy (TUE) score, which combines sub-metrics evaluating how correctly and efficiently an agent uses external tools (including proper tool selection and correct parameter usage in tool calls) into one measure. By condensing multiple criteria, composite metrics enable high-level comparison and benchmarking of Agentic AI systems. For instance, AgentBench [168] is a comprehensive benchmark that evaluates agents across a diverse range of tasks and environments (from operating system manipulation to web shopping), effectively providing a composite performance profile of an agent. Such aggregated scores highlight if an agent performs strongly across the board or if it excels in some dimensions while underperforming in others. It is important to interpret composite scores in light of their components: a single number can mask specific weaknesses (e.g., an agent might achieve a high overall score by doing well in task completion and coordination, yet still have poor explainability). Therefore, composite metrics are most informative when accompanied by a breakdown of the agent’s performance per category.
@@ 8 Security and Privacy in LLM-based Agentic Multi-Agent Architectures
In this section, we discuss the security and privacy in AMAS.
Table 8: Governance Dimensions for Agentic AI Systems
Governance Component Description Example Tools or Methods
Regulatory Compliance Aligning with legal and industry standards for responsible AI [169] NIST AI RMF [87], EU AI Act [170], GDPR [91], HIPAA [92]
Auditability and Logging Capturing traceable records of agent actions and decisions [171] Decision provenance, immutable logs, blockchain-based traceability
Policy Enforcement Enforcing system-level operational, ethical, and security constraints [53, 172] RBAC/ABAC policies [173], formal logic rules, dynamic sandboxing [60, 174, 175]
Human Oversight Human-in-the-loop [151] decision checkpoints and override capabilities Human review agents, interactive dashboards, compliance checkpoints
Risk Monitoring Continuous identification and mitigation of systemic or emergent risks [176] Model risk scanners, out-of-distribution detectors, reinforcement learning audit modules
Explainability Governance Ensuring interpretability standards are met in decision pipelines [177] LIME/SHAP [178, 179, 180], decision trees [181, 182], interpretable surrogate models [183, 184, 185]
Adaptive Governance Updating governance mechanisms based on model evolution and deployment context [144, 146] Governance-as-code [114, 186], auto-updated rule engines [187, 188, 189], learning-based policy adaptation [126, 190]
@@ 8.1 Security Mechanisms
Agentic AI systems, composed of loosely coupled yet collaboratively functioning LLM-based agents, introduce an expanded attack surface relative to conventional AI agents [191]. Ensuring the security of such systems necessitates a multi-layered defense architecture that addresses data protection, execution integrity, inter-agent communication, and model robustness [192]. Among the foundational techniques employed are encryption [193, 194], access control [195, 171], adversarial defense [196, 197], and runtime monitoring[198, 199] each adapted to the unique demands of decentralized multi-agent environments.
@@
Encryption. plays an important role in safeguarding data exchanged between agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved [194, 61, 200]. Agentic workflows often include inter-agent handoff of partially processed results, models, or prompts. Implementations such as SSL/TLS, homomorphic encryption [201, 200], and secure enclaves (e.g., Intel SGX) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
@@
Access control. becomes crucial when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities. For instance, in systems like AutoGen[18] and CrewAI [81] where agents take on specialized roles (e.g., summarizer, planner, coder) [202], enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [18, 198, 203]. Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [204, 205] and Attribute-Based Access Control (ABAC) [206, 205] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels [5].
@@
Adversarial robustness. is a growing concern as LLM-based agents are susceptible to prompt injection, manipulation through poisoned tool outputs, or coordination disruption via malformed intermediate results [207, 208]. Recent studies have shown that multi-agent LLM frameworks can be destabilized by adversarially crafted outputs from one compromised agent propagating misleading information to others [208, 192, 115, 55, 116, 209]. Adversarial training methods, such as input perturbation, reward shaping, and contrastive learning, can partially mitigate these vulnerabilities [210, 115]. Integrating safety constraints and verifying tool responses before execution are also effective mitigation strategies.
@@
Runtime monitoring. systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity [211]. Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers [212, 213, 125]. For example, Microsoft’s Copilot governance layers monitor anomalous agent behavior across sessions to ensure compliant execution and flag potentially harmful interactions.
@@
As Agentic AI systems continue to expand into mission-critical domains, developing standardized, scalable security mechanisms will be paramount. Future approaches must include zero-trust frameworks, secure multi-party computation, and formal verification of inter-agent protocols to ensure safe and resilient operation across decentralized autonomous agent collectives.
@@ 8.2 Privacy-Preserving Techniques
Differential privacy (DP). DP offers mathematically grounded guarantees by injecting statistical noise into outputs, ensuring that individual user contributions cannot be re-identified. In multi-agent LLM systems, DP can be applied during training or at inference-time when agents exchange information. For example, Google’s implementation of DP in federated learning frameworks can be extended to distributed agentic systems, where agents collaboratively train or fine-tune local models without exposing raw data [214, 215]. DP-SGD and privacy budgets ($\epsilon$-differentials) can regulate information exposure during policy updates or collaborative planning in real-time decision-making agents.
Data minimization. Agentic AI systems can mitigate exposure risks by limiting the scope, granularity, and duration of data collected or retained during task execution. For instance, temporary memory buffers used in systems like ChatDev or ReAct-based pipelines are cleared once subgoals are completed, preventing persistent storage of unnecessary user data. Furthermore, anonymization and pseudony [215] techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
Secure computation. Techniques including secure multi-party computation (SMPC), homomorphic encryption, and trusted execution environments (TEEs) enable agents to perform computations over encrypted or obfuscated data without compromising privacy. In scenarios where agents collaborate across different organizational boundaries (e.g., federated medical agents or cross-silo industrial agents), SMPC allows joint computations such as diagnostics or anomaly detection without data leakage. Homomorphic encryption, while computationally expensive, is increasingly being explored to allow arithmetic operations on encrypted vectors used in RAG workflows.
@@ 8.3 Compliance and Governance
As Agentic AI systems grow in capability and autonomy, ensuring regulatory compliance and instituting robust governance mechanisms become imperative. Unlike traditional AI agents, Agentic AI systems operate with greater autonomy, persistent memory, and complex decision flows necessitating layered oversight to manage legal, ethical, and societal implications. Effective governance in this context spans three critical dimensions: adherence to regulatory standards, system-level auditability, and enforceable policy frameworks.
@@
Regulatory standards. Standards for regulation provide the baseline requirements that all AI systems, including agentic architectures, must meet. Frameworks such as the NIST AI Risk Management Framework (AI RMF) and the EU AI Act define principles for trustworthy AI including transparency, accountability, and fairness. These standards are especially relevant for LLM-based AMAS that interact in high-stakes domains like healthcare, finance, defense, or transportation. For example, the EU AI Act classifies certain autonomous systems as “high-risk,” requiring continuous risk monitoring of such systems.
@@
Auditability. This is critical in ensuring transparency and facilitating post-hoc accountability. Each decision, plan, or interaction within an agentic system should be logged with timestamps, context, agent role, and justification. Techniques such as decision provenance and action traceability enable this, allowing regulators or internal auditors to reconstruct how decisions were reached. For instance, in systems like AutoGen or MetaGPT, where agents assume specialized roles (e.g., researcher, coder, reviewer), audit trails can capture role-specific actions and flag inconsistencies, bias amplification, or security violations. Blockchain-based audit logs are also being explored to ensure immutability and verifiability of multi-agent interactions.
@@
Policy enforcement. Policy governs what Agentic AI system can and cannot do, and under which conditions. These policies must be codified into the orchestration layers or meta-agent governance modules that manage agent interactions. Examples include enforcing memory expiration policies to avoid data retention violations or restricting access to external tools based on role and authentication level. Role-based access control (RBAC) and attribute-based access control (ABAC) are essential for enforcing differentiated privileges across agent sub-components. Furthermore, real-time monitoring systems can halt or flag agent activity that deviates from pre-specified ethical constraints or operational bounds, using formal verification tools like TLA+ or symbolic execution engines.
@@
Emerging best practices also include the creation of AI governance boards, the adoption of governance-as-code platforms, and the integration of adaptive governance layers that evolve as the Agentic AI systems scale or change context. These practices are designed to meet not only current standards like ISO/IEC 42001 for AI management systems, but also prepare for future regulatory evolutions. Below, we summarize the core governance components required for secure and compliant Agentic AI deployment in Table 8.